Mast Corporation seeks your assistance in developing cash and other budget information for May, June, and July. At April 30, the company had cash of $11,000, accounts receivable of $876,000, inventories of $128,520, and accounts payable of $36,202. The budget is to be based on the following assumptions.

  • Each month’s sales are billed on the last day of the month.
  • Customers are allowed a 2 percent discount if payment is made within 10 days after the billing date. Receivables are recorded in the accounts at their gross amounts (not net of discounts).
  • The billings are collected as follows: 65 percent within the discount period, 20 percent by the end of the month, and 12 percent by the end of the following month. Three percent is uncollectible.

Purchase data are as follows.

  • Of all purchases of merchandise and selling, general, and administrative expenses, 56 percent is paid in the month purchased and the remainder in the following month.
  • The number of units in each month’s ending inventory equals 120 percent of the next month’s units of sales.
  • The cost of each unit of inventory is $9.
  • Selling, general, and administrative expenses, of which $4,000 is depreciation, equal 15 percent of the current month’s sales.
  • Actual and projected sales follow:
Dollars Units
March $ 165,200 11,800
April 183,400 13,100
May 166,600 11,900
June 159,600 11,400
July 168,000 12,000
August 14,800 12,200


Required:

a. Compute the budgeted purchases in dollars for May.
b. Compute the budgeted purchases in dollars for June.
c. Compute the budgeted cash collections during May. (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)
d. Compute the budgeted cash disbursements during June. (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)
e. Compute the budgeted number of units of inventory to be purchased during July.
Answer #1

Above question is about to find out budgeted purchases in dollars for the months of may ,June and cash collections during may and cash disbursement during June and budgeted number of units purchased during July. Refer below images for more detailed solution.solution : a) Budgeted purchases in dollars for may BB+ P = Sales & EB (120-4 X11900) + p = 1900+ (1207419400) 14280 tp = 119c) Budgeted cash collections during May: $ 166600 x 65.10 x 98% = $106124.2 $ 166600 x 20% - $ 333200 $ 183400 X 120. $ 22008Note * ** points : from part a of this problem. from part b. of this problem. Given 56% payments are done in the Remaining 44